# From Unimodal to Multimodal: Scaling up Projectors to Align Modalities

### v1

@@ A.2 CKA vs Graph structure
A.3 Comparison to LiLT
Tables A.1 and A.2 report the zero-shot domain classification and retrieval performance of LiLT models Khan & Fu (2023). The vision encoder is initialized with the DeiT base model Touvron et al. (2021), and the text encoder is from SimCSE Gao et al. (2021). The LilTDA-base model is trained by duplicating and appending the last transformer layer, while only unlocking the last encoder and projector layers. The LilTLwA-base model introduces trainable layerwise adapters for both the vision and text encoders. LiLT public checkpoints are trained on 500k image-caption pairs from the COCO dataset. However, LiLT’s performance lags behind CLIP models and our DINOv2-ARL projector model, primarily due to suboptimal encoder pairs and limited concept coverage in the COCO training set for alignment.
@@ A.3 Comparison to LiLT
A.4 Encoder Pairs Ablations
@@
A naive approach to choosing the best encoder pair is to chose the unimodal encoders with highest performance in their respective modalities, but it’s not straightforward which benchmarks can be more predictive of ease of alignment. To demonstrate this, we consider the same ablation as above, but with DINOv2 and 14 different text encoders from the SentenceTransformers Reimers & Gurevych (2019) library. We consider 2 types of text model benchmarks. 1. Sentence Embedding task or Semantic Textual Similarity (STS) is the task of evaluating how similar two texts are in terms of meaning. These models take a source sentence and a list of sentences and return a list of similarity scores. The task is evaluated using Spearman’s Rank Correlation. We average over 14 datasets reported in Reimers & Gurevych (2019; 2024). 2. Semantic Search (SS) is the task of retrieving relevant documents or passages based on the semantic content of a query. Rather than relying solely on keyword matching, semantic search models generate embeddings for both the query and the documents, allowing for retrieval based on contextual and conceptual similarity and is evaluated using Normalized Discounted Cumulative Gain (nDCG), which measure the relevance of retrieved documents in ranked lists. We average over 6 datasets reported in Reimers & Gurevych (2019; 2024).
@@
A.5 Data Curation Implementation Details
We streamline our class collection process by precomputing CLIP text embeddings for LAION-400M and CLIP image prototype embeddings for various concepts, allowing us to run different collection methods without needing to recompute embeddings. The embedding process takes just 12 hours on two nodes with 4 A6000 GPUs each. Class-level collection is performed using GPU-accelerated PyTorch code on a single GPU, completing in under an hour. While image-to-image-prototype collection, as in Oquab et al. (2023), could yield higher-quality results, it demands significantly more GPU resources due to the need to create CLIP embeddings for all LAION-400M images. We find that caption-image-concept similarity performs well for image classification accuracy. To support efficient multi-modal model training, we release the LAION-CLASS-Collected parquets for research use.
A.6 Projector training details
@@
A.7 Multi-Lingual Full Results
Another significant advantage of using only Projectors to align modalities is the ability to swap the text encoder with multi-lingual encoders trained on various languages, thus potentially extending a CLIP model to accommodate any language. This feature is particularly beneficial for low-resource languages. We demonstrate the feasibility of this approach by training projectors to align the DINOv2 visual encoder with the paraphrase-multilingual-v2 text encoder, using a dataset consisting solely of English image-caption pairs. We selected this specific text encoder as it showed the highest compatibility in terms of CKA with DINOv2. Subsequently, we evaluated the performance of our model on multi-lingual image retrieval using the XTD dataset Aggarwal & Kale (2020a) and on multi-lingual image classification using the ImageNet dataset. For multi-lingual classification, we translate our VDT prompts Maniparambil et al. (2023) to the languages being considered using the nllb-700M model Costa-jussà et al. (2022) and then use the same prompts for all the models being considered including ours.
For both multi-lingual classification and retrieval tasks, our comparisons are structured into two categories as delineated in Table A.4 and Table A.3. The lower sections of each of these tables list models trained exclusively with English captions, more specifically the CLIP-VIT-L models from OpenAI and LAION trained on 400 million image caption pairs of WIT dataset and LAION400M dataset respectively. The upper sections of these tables feature models trained with translated captions, including those employing contrastive training with multi-lingual image-caption pairs such as CLIP-models based on the LAION5B multi-lingual dataset, which contains image-caption pairs in over 100 languages. We also compare against, M-CLIP Chen et al. (2023a) models that are trained using English and translated captions to align a multi-lingual text encoder with CLIP’s original text encoder through contrastive learning, thereby enhancing performance on multi-lingual tasks. Additionally we also compare against the NLLB-CLIP Visheratin (2023) models developed through LiT Zhai et al. (2022) techniques, coupling a frozen CLIP visual encoder with an unfrozen multi-lingual text encoder using translated captions from the smaller LAION-COCO dataset. We compare against only model sizes of up to ViT-Large for fair comparison.
@@ A.8 Dataset scale
A.9 sDCI benchmark results
We evaluate our method on the Densely Captioned Images (DCI) dataset Urbanek et al. (2024), which contains 7,805 images with mask-aligned descriptions averaging over 1,000 words each. To accommodate current models’ token limits, the authors also provide sDCI, a summarized version with CLIP-compatible 77-token captions generated by LLMs.
@@
A.10 0-Shot Classification and Retrieval Evaluation Datasets
@@
ImageNet Deng et al. (2009): A large-scale dataset with 1000 object categories, widely used as a benchmark for image classification tasks. It contains over 1.2 million training images and 50,000 validation images, with each image labeled with one of 1000 object classes.
@@
ImageNetV2 Recht et al. (2019): A newer version of ImageNet designed to test the robustness of models trained on the original ImageNet. It features 10,000 new test images collected using the same procedure as the original, but addressing certain biases in the original dataset.
@@
Caltech101 Li et al. (2022): A dataset containing pictures of objects belonging to 101 categories, plus a background category. It includes about 40 to 800 images per category, with most categories having about 50 images. The dataset is known for its high intra-class variability.
@@
Oxford-IIIT Pet Parkhi et al. (2012): A 37-category pet dataset with roughly 200 images for each class, featuring different breeds of cats and dogs. It includes pixel-level trimap segmentations and breed-level labels for each image.
@@
Stanford Cars Krause et al. (2013): A dataset of 196 car classes, totaling 16,185 images. Classes are at the level of Make, Model, Year (e.g., 2012 Tesla Model S). It includes 8,144 training images and 8,041 testing images, with bounding box annotations.
@@
Oxford Flowers102 Nilsback & Zisserman (2008): A 102 category dataset consisting of 102 flower categories common to the UK. It contains 40 to 258 images per class and provides segmentation data for each image. The dataset is particularly challenging due to the fine-grained nature of the categories.
@@
Food101 Bossard et al. (2014): A large dataset of 101 food categories, with 101,000 images. It features 1000 images per food class, with 250 test images and 750 training images per class. The training images are not manually cleaned, adding a level of noise to the dataset.
@@
FGVC Aircraft Maji et al. (2013): A fine-grained visual classification dataset with 10,200 images of aircraft, spanning 100 aircraft models. Each model is associated with a specific variant, manufacturer, family, and collection. The dataset includes 6,667 training images and 3,333 test images.
@@
SUN397 Rouach et al. (2020): A scene recognition dataset with 397 categories and 108,754 images, covering a large variety of environmental scenes under various lighting conditions. It provides at least 100 images per class and has been used extensively for scene recognition tasks.
@@
Caltech-UCSD Birds-200-2011 (CUB) Wah et al. (2011): A dataset for fine-grained image classification with 200 bird species, containing 11,788 images. Each image has detailed annotations including 15 part locations, 312 binary attributes, and 1 bounding box. It’s widely used for fine-grained visual categorization research.
@@
UCF101 Soomro et al. (2012): An action recognition dataset with 101 action categories, consisting of realistic action videos collected from YouTube. It contains 13,320 videos from 101 action categories, with videos exhibiting large variations in camera motion, object appearance and pose, illumination conditions, and more.
@@
Flickr30k Plummer et al. (2015): A dataset containing 31,783 images collected from Flickr, each paired with 5 crowd-sourced captions. It focuses on describing the objects and actions in everyday scenes. The dataset is split into 29,783 training images, 1000 validation images, and 1000 test images.
COCO Lin et al. (2014): A large-scale dataset for object detection, segmentation, and captioning, which we use for its image-caption pairs in the retrieval task. It features over 330,000 images, each with 5 captions. The dataset includes 80 object categories and instance segmentation masks, making it versatile for various computer vision tasks.
These datasets comprehensively evaluate a model’s ability to perform zero-shot classification across various domains and its capacity for cross-modal retrieval. By using this diverse set of benchmarks, we can assess the generalization capabilities of our approach compared to existing CLIP models. We use Visually Descriptive Class-Wise prompts from Maniparambil et al. (2023) to enable the unimodal-text encoder in our DINOv2-ARL projector model to better identify the zero-shot classes of the downstream datasets.
A.10.1 Concept Coverage Collection datasets
@@
BirdSnap Berg et al. (2014): A fine-grained dataset consisting of 49,829 images of 500 North American bird species. The images are annotated with species labels, and the dataset is primarily used for species classification and fine-grained recognition tasks.
@@
Caltech101 Li et al. (2022): A dataset containing pictures of objects belonging to 101 categories, plus a background category. It includes about 40 to 800 images per category, with most categories having about 50 images. The dataset is known for its high intra-class variability.
@@
EuroSAT Helber et al. (2019): A satellite image dataset with 10 categories related to land use classification (e.g., forests, rivers, residential areas). It contains 27,000 labeled images, with 2700 images per class, widely used in remote sensing and geospatial tasks.
@@
FGVC Aircraft Maji et al. (2013): A fine-grained classification dataset with 10,000 images of 100 aircraft model variants from 70 manufacturers. It is used for distinguishing between visually similar objects in fine-grained recognition tasks.
@@
Flowers102 Nilsback & Zisserman (2008): A dataset containing 102 flower categories, commonly used for fine-grained classification tasks. It has a total of 8,189 images, with 40 to 258 images per category, and is organized into a training, validation, and test set.
@@
Food101 Bossard et al. (2014): A dataset containing 101,000 images of 101 food categories. Each category has 750 training images and 250 test images, commonly used for food classification and recognition tasks.
@@
GTSRB Stallkamp et al. (2012): The German Traffic Sign Recognition Benchmark dataset, containing over 50,000 images of 43 different traffic sign classes. It is designed for multi-class classification tasks in the context of traffic sign recognition.
@@
ImageNet Deng et al. (2009): A large-scale dataset with 1,000 object categories, widely used as a benchmark for image classification tasks. It contains over 1.2 million training images and 50,000 validation images, with each image labeled with one of 1,000 object classes.
@@
Oxford Pets Parkhi et al. (2012): A dataset of 7,349 images, containing 37 categories of pets (both cats and dogs). Each image is annotated with species and breed information, commonly used for image classification and segmentation tasks.
@@
RESISC45 Cheng et al. (2017): A dataset of remote sensing images used for scene classification, containing 31,500 images across 45 scene classes. Each class has 700 images with variations in resolution, scale, and orientation.
@@
Stanford Cars Krause et al. (2013): A dataset with 16,185 images of 196 car models, annotated by make, model, and year. The dataset is designed for fine-grained classification and recognition tasks of vehicles.
@@
Pascal VOC 2007 Everingham et al. (2015): A dataset for object detection, segmentation, and classification, containing 9,963 images of 20 object categories. It is widely used for benchmarking models in computer vision tasks.
@@
SUN397 Rouach et al. (2020): A large-scale scene understanding dataset with 397 categories and 108,754 images. It covers a wide range of environments, from natural to man-made scenes, commonly used for scene classification tasks.
@@
UCF101 Soomro et al. (2012): A video dataset consisting of 13,320 videos across 101 human action categories. It is widely used for action recognition tasks in video analysis and computer vision research.

### v2

@@ A.5 Embedding Graph structures visualized
A.6 Comparison to LiLT
Tables A.4 and A.5 report the zero-shot domain classification and retrieval performance of LiLT models [23]. The vision encoder is initialized with the DeiT base model [56], and the text encoder is from SimCSE [17]. The LilTDA-base model is trained by duplicating and appending the last transformer layer, while only unlocking the last encoder and projector layers. The LilTLwA-base model introduces trainable layerwise adapters for both the vision and text encoders. LiLT public checkpoints are trained on 500k image-caption pairs from the COCO dataset. However, LiLT’s performance lags behind CLIP models and our DINOv2-ARL projector model, primarily due to suboptimal encoder pairs and limited concept coverage in the COCO training set for alignment.
@@ A.6 Comparison to LiLT
A.7 Encoder Pairs Ablations
@@
A naive approach to choosing the best encoder pair is to chose the unimodal encoders with highest performance in their respective modalities, but it’s not straightforward which benchmarks can be more predictive of ease of alignment. To demonstrate this, we consider the same ablation as above, but with DINOv2 and 14 different text encoders from the SentenceTransformers [47] library. We consider 2 types of text model benchmarks. 1. Sentence Embedding task or Semantic Textual Similarity (STS) is the task of evaluating how similar two texts are in terms of meaning. These models take a source sentence and a list of sentences and return a list of similarity scores. The task is evaluated using Spearman’s Rank Correlation. We average over 14 datasets reported in [47, 48]. 2. Semantic Search (SS) is the task of retrieving relevant documents or passages based on the semantic content of a query. Rather than relying solely on keyword matching, semantic search models generate embeddings for both the query and the documents, allowing for retrieval based on contextual and conceptual similarity and is evaluated using Normalized Discounted Cumulative Gain (nDCG), which measure the relevance of retrieved documents in ranked lists. We average over 6 datasets reported in [47, 48].
@@
A.8 Data Curation Implementation Details
We streamline our class collection process by precomputing CLIP text embeddings for LAION-400M and CLIP image prototype embeddings for various concepts, allowing us to run different collection methods without needing to recompute embeddings. The embedding process takes just 12 hours on two nodes with 4 A6000 GPUs each. Class-level collection is performed using GPU-accelerated PyTorch code on a single GPU, completing in under an hour. While image-to-image-prototype collection, as in [41], could yield higher-quality results, it demands significantly more GPU resources due to the need to create CLIP embeddings for all LAION-400M images. We find that caption-image-concept similarity performs well for image classification accuracy. To support efficient multi-modal model training, we release the LAION-CLASS-Collected parquets for research use.
A.9 Projector training details
@@
A.10 0-shot Segmentation Evaluation
In DINOV2-ARL, we perform 0-shot segmentation by computing cosine similarities between each patch and all the ground truth classes and subsequently upscaling to the target size. Each patch is then classified into a corresponding class. Consistent with previous studies, the intersection over union (IoU) is computed solely for the foreground classes. In the zero-shot segmentation process of CLIP models, we employ a technique similar to [64] to alleviate the opposite visualization problem in CLIP models [28]. The patch embeddings from the penultimate layer are passed through the value layer and output MLP of the final self-attention block, followed by projection into the joint embedding space using the vision projector. Meanwhile, our DINOv2-ARL model considers patch embeddings projected into the joint embedding space by the patch projector and augments them with the projected CLS token in a residual manner.
A.11 Multi-Lingual Full Results
Another significant advantage of using only Projectors to align modalities is the ability to swap the text encoder with multi-lingual encoders trained on various languages, thus potentially extending a CLIP model to accommodate any language. This feature is particularly beneficial for low-resource languages. We demonstrate the feasibility of this approach by training projectors to align the DINOv2 visual encoder with the paraphrase-multilingual-v2 text encoder, using a dataset consisting solely of English image-caption pairs. We selected this specific text encoder as it showed the highest compatibility in terms of CKA with DINOv2. Subsequently, we evaluated the performance of our model on multi-lingual image retrieval using the XTD dataset [2] and on multi-lingual image classification using the ImageNet dataset. For multi-lingual classification, we translate our VDT prompts [33] to the languages being considered using the nllb-700M model [12] and then use the same prompts for all the models being considered including ours.
For both multi-lingual classification and retrieval tasks, our comparisons are structured into two categories as delineated in Table A.7 and Table A.6. The lower sections of each of these tables list models trained exclusively with English captions, more specifically the CLIP-VIT-L models from OpenAI and LAION trained on 400 million image caption pairs of WIT dataset and LAION400M dataset respectively. The upper sections of these tables feature models trained with translated captions, including those employing contrastive training with multi-lingual image-caption pairs such as CLIP-models based on the LAION5B multi-lingual dataset, which contains image-caption pairs in over 100 languages. We also compare against, M-CLIP [9] models that are trained using English and translated captions to align a multi-lingual text encoder with CLIP’s original text encoder through contrastive learning, thereby enhancing performance on multi-lingual tasks. Additionally we also compare against the NLLB-CLIP [58] models developed through LiT [63] techniques, coupling a frozen CLIP visual encoder with an unfrozen multi-lingual text encoder using translated captions from the smaller LAION-COCO dataset. We compare against only model sizes of up to ViT-Large for fair comparison.
@@ A.12 Dataset scale
A.13 sDCI benchmark results
We evaluate our method on the Densely Captioned Images (DCI) dataset [57], which contains 7,805 images with mask-aligned descriptions averaging over 1,000 words each. To accommodate current models’ token limits, the authors also provide sDCI, a summarized version with CLIP-compatible 77-token captions generated by LLMs.
@@
A.14 0-Shot Classification and Retrieval Evaluation Datasets
@@
ImageNet [13]: A large-scale dataset with 1000 object categories, widely used as a benchmark for image classification tasks. It contains over 1.2 million training images and 50,000 validation images, with each image labeled with one of 1000 object classes.
@@
ImageNetV2 [46]: A newer version of ImageNet designed to test the robustness of models trained on the original ImageNet. It features 10,000 new test images collected using the same procedure as the original, but addressing certain biases in the original dataset.
@@
Caltech101 [27]: A dataset containing pictures of objects belonging to 101 categories, plus a background category. It includes about 40 to 800 images per category, with most categories having about 50 images. The dataset is known for its high intra-class variability.
@@
Oxford-IIIT Pet [43]: A 37-category pet dataset with roughly 200 images for each class, featuring different breeds of cats and dogs. It includes pixel-level trimap segmentations and breed-level labels for each image.
@@
Stanford Cars [25]: A dataset of 196 car classes, totaling 16,185 images. Classes are at the level of Make, Model, Year (e.g., 2012 Tesla Model S). It includes 8,144 training images and 8,041 testing images, with bounding box annotations.
@@
Oxford Flowers102 [39]: A 102 category dataset consisting of 102 flower categories common to the UK. It contains 40 to 258 images per class and provides segmentation data for each image. The dataset is particularly challenging due to the fine-grained nature of the categories.
@@
Food101 [6]: A large dataset of 101 food categories, with 101,000 images. It features 1000 images per food class, with 250 test images and 750 training images per class. The training images are not manually cleaned, adding a level of noise to the dataset.
@@
FGVC Aircraft [31]: A fine-grained visual classification dataset with 10,200 images of aircraft, spanning 100 aircraft models. Each model is associated with a specific variant, manufacturer, family, and collection. The dataset includes 6,667 training images and 3,333 test images.
@@
SUN397 [49]: A scene recognition dataset with 397 categories and 108,754 images, covering a large variety of environmental scenes under various lighting conditions. It provides at least 100 images per class and has been used extensively for scene recognition tasks.
@@
Caltech-UCSD Birds-200-2011 (CUB) [59]: A dataset for fine-grained image classification with 200 bird species, containing 11,788 images. Each image has detailed annotations including 15 part locations, 312 binary attributes, and 1 bounding box. It’s widely used for fine-grained visual categorization research.
@@
UCF101 [53]: An action recognition dataset with 101 action categories, consisting of realistic action videos collected from YouTube. It contains 13,320 videos from 101 action categories, with videos exhibiting large variations in camera motion, object appearance and pose, illumination conditions, and more.
@@
Flickr30k [44]: A dataset containing 31,783 images collected from Flickr, each paired with 5 crowd-sourced captions. It focuses on describing the objects and actions in everyday scenes. The dataset is split into 29,783 training images, 1000 validation images, and 1000 test images.
COCO [29]: A large-scale dataset for object detection, segmentation, and captioning, which we use for its image-caption pairs in the retrieval task. It features over 330,000 images, each with 5 captions. The dataset includes 80 object categories and instance segmentation masks, making it versatile for various computer vision tasks.
These datasets comprehensively evaluate a model’s ability to perform zero-shot classification across various domains and its capacity for cross-modal retrieval. By using this diverse set of benchmarks, we can assess the generalization capabilities of our approach compared to existing CLIP models. We use Visually Descriptive Class-Wise prompts from [33] to enable the unimodal-text encoder in our DINOv2-ARL projector model to better identify the zero-shot classes of the downstream datasets.
A.14.1 Concept Coverage Collection datasets
@@
BirdSnap [4]: A fine-grained dataset consisting of 49,829 images of 500 North American bird species. The images are annotated with species labels, and the dataset is primarily used for species classification and fine-grained recognition tasks.
@@
Caltech101 [27]: A dataset containing pictures of objects belonging to 101 categories, plus a background category. It includes about 40 to 800 images per category, with most categories having about 50 images. The dataset is known for its high intra-class variability.
@@
EuroSAT [19]: A satellite image dataset with 10 categories related to land use classification (e.g., forests, rivers, residential areas). It contains 27,000 labeled images, with 2700 images per class, widely used in remote sensing and geospatial tasks.
@@
FGVC Aircraft [31]: A fine-grained classification dataset with 10,000 images of 100 aircraft model variants from 70 manufacturers. It is used for distinguishing between visually similar objects in fine-grained recognition tasks.
@@
Flowers102 [39]: A dataset containing 102 flower categories, commonly used for fine-grained classification tasks. It has a total of 8,189 images, with 40 to 258 images per category, and is organized into a training, validation, and test set.
@@
Food101 [6]: A dataset containing 101,000 images of 101 food categories. Each category has 750 training images and 250 test images, commonly used for food classification and recognition tasks.
@@
GTSRB [54]: The German Traffic Sign Recognition Benchmark dataset, containing over 50,000 images of 43 different traffic sign classes. It is designed for multi-class classification tasks in the context of traffic sign recognition.
@@
ImageNet [13]: A large-scale dataset with 1,000 object categories, widely used as a benchmark for image classification tasks. It contains over 1.2 million training images and 50,000 validation images, with each image labeled with one of 1,000 object classes.
@@
Oxford Pets [43]: A dataset of 7,349 images, containing 37 categories of pets (both cats and dogs). Each image is annotated with species and breed information, commonly used for image classification and segmentation tasks.
@@
RESISC45 [11]: A dataset of remote sensing images used for scene classification, containing 31,500 images across 45 scene classes. Each class has 700 images with variations in resolution, scale, and orientation.
@@
Stanford Cars [25]: A dataset with 16,185 images of 196 car models, annotated by make, model, and year. The dataset is designed for fine-grained classification and recognition tasks of vehicles.
@@
Pascal VOC 2007 [14]: A dataset for object detection, segmentation, and classification, containing 9,963 images of 20 object categories. It is widely used for benchmarking models in computer vision tasks.
@@
SUN397 [49]: A large-scale scene understanding dataset with 397 categories and 108,754 images. It covers a wide range of environments, from natural to man-made scenes, commonly used for scene classification tasks.
@@
UCF101 [53]: A video dataset consisting of 13,320 videos across 101 human action categories. It is widely used for action recognition tasks in video analysis and computer vision research.